OpenStudy (anonymous):

how to simplify the square roo tof 54

OpenStudy (anonymous):

Find a factor of 54 that's a square number: 1 - Yes but ignore it. 4 - No. 9 - Yes. Now split 54 into two factors and separate the surd: sqrt(54)=sqrt(9*6)=sqrt(9)*sqrt(6)=3sqrt(6)

OpenStudy (anonymous):

7.348

OpenStudy (anonymous):

what

OpenStudy (anonymous):

No, don't work it out. Your answer is 3sqrt(6).